OpenStudy (anonymous):

can anyone solve this ? Factor the trinomial. a^2 − 3a − 40

OpenStudy (anonymous):

Find two numbers that multiply out to AC and add up to B. Thus, two numbers that multiply to -40 and add up to -3.

OpenStudy (anonymous):

Those values are -8 and 5. Thus, the factored form is (a-5)(a+8)=0

OpenStudy (anonymous):

(a+5)(a-8)=0 (a+5)=0 and (a-8)=0 solve a

OpenStudy (anonymous):

what about r^2 − 7r − 10?